Find the values of a for which the function f (x) = cos x - sin x + ax + b is decreasing.

Grade:12

1 Answers

Arun Kumar IIT Delhi
9 years ago
Hello Student,
f'(x)=-sinx-cox+a
Min(-sinx-cosx)= -root(2)=-1.4

f'(x)>a-1.4
f'(x)>0
a-1.4>=0
a in (1.4, inf)
Thanks & Regards
Arun Kumar
Btech, IIT Delhi
Askiitians Faculty